OpenStudy (anonymous):

Solve the following equation by transforming it into a perfect square trinomial. x2 – 4x = 5 A. {–5, 13} B. {–7, 11} C. {1, 7} D. {–1, 5}

OpenStudy (anonymous):

so when creating a perfect square trinomial, we add a number to both sides so we can factor the equation easier and we find this number we add using (b/2)^2 b=-4 in this case so (-4/2)^2=(-2)^2=4 so now we add 4 to both sides x^2–4x+4=5+4 this makes x^2-4x+4 into a perfect square of (x-2)^2 so (x-2)^2=9 now to solve for x we square root the equation to get rid of the exponent x-2=+-3 and now we set x-2 equal to both +3 and -3 since both squared =9 x-2=-3 x-2=+3 for both you add 2 to both sides x=-1 x=5 so the answer is D